OpenStudy (mathhelp346):

The sum of the digits of a two-digit number is 8. If 16 is added to the original number, the result is 3 times the original number with its digits reversed. Find the original number. I got: {t+u=8 10t+u=3(10u+t)+16 But I don't know how to solve it? And is it right?

OpenStudy (anonymous):

62

OpenStudy (mathhelp346):

how did you solve it

OpenStudy (anonymous):

take this number as ab, so a+b=8 and (ab)+16=3(ba) which means 10a+b+16 =3(10b+a) If you solve the equation, you get 29b-7a=16 adn you know that a+b=8, if you solve them both, you get 36b=72 so b=2 and a=6 ab=62

OpenStudy (mathhelp346):

wait what?

OpenStudy (mathhelp346):

how is it (ab)+16=3(ba)?

OpenStudy (anonymous):

The mistake you make is this: You should think that if you add 16 to the number, you get 3 times of this so 16 should be on the leftside

OpenStudy (mathhelp346):

shouldn't it be 3(ab)+16

OpenStudy (anonymous):

got it_?

OpenStudy (mathhelp346):

no

OpenStudy (mathhelp346):

how is (ab)+16=3(ba)?

OpenStudy (anonymous):

The problem says " If 16 is added to the original number, the result is 3 times the original number with its digits reversed" which means if you add 16 to (ab), you get 3 times of(ba)

OpenStudy (mathhelp346):

oh ok

OpenStudy (mathhelp346):

i don't know you got 62

OpenStudy (anonymous):

is it wrong?

OpenStudy (mathhelp346):

when i solved 10a+b+16 =3(10b+a) I got 10a+16=29b+3a

OpenStudy (mathhelp346):

no wait i got 7a+16=29b and i don't know what to do from there

OpenStudy (anonymous):

10a+b+16=30b+3a so b+16=30b-7a and 29b-7a=16

OpenStudy (mathhelp346):

shouldn't u subtract 3t though

OpenStudy (anonymous):

ok now you geot the true equation 7a+16=29b... you know that a+b=8, for it is in the question.Now write these equations in two lines and solve them both

OpenStudy (anonymous):

ok forget a and b..let's do it your way.. let the number be (ut).. you know that u+t=8 and 10u+t+16=3(10t+u)

OpenStudy (anonymous):

10u+t+16=30t+3u so 7u-29t=-16 so 29t-7u=16

OpenStudy (mathhelp346):

yea..

OpenStudy (anonymous):

29t-7u=16 and u+t=8 so u=8-t. if you put 8-t in the first equation; 29t-7(8-t)=16 so 29t-56+7t=16 and t=2 here so u=6

OpenStudy (mathhelp346):

so if t=2 and u=6, how do you know to put it 62? like how do you know to not put it as 26?

OpenStudy (anonymous):

because in the beginning you named the number as (ut) which means u=6 and t=2 and ut is 62.. we solved it according to this information we put 10u instead of u so it is 62

OpenStudy (mathhelp346):

o ok thanks so much

OpenStudy (anonymous):

ur welcome:)